kira:

7. How many grams of sodium chloride are in 250 mL of a 2.5M NaCl solution?

cpl106:

Given, volume of solution is 250ml or 0.25 liter. Molarity of NaCl = 0.50M Now, according to molarity formula M = n/V(in L) where , n = no. of mole of solute(NaCl) put the value 0.5 = n/0.25 => n = 0.5×0.25 => n = 0.125mole So, number of mole of solute (NaCl) is 0.125mole present in solution . Now , molecular mass of NaCl is NaCl = 23+ 35.5 = 58.5 Now, according to formula number if mole = mass of NaCl/molecular mass => 0.125 = mass of NaCl/58.5 => mass of NaCl = 0.125×58.5 = 7.3125g Hence, 7.3125 grams of sodium chloride is present in 250mL.